Kathy is 6yrs older than her cat. Next year she will be twice as old as her cat will be. How old is her cat now?
Answer by oberobic(2304)

Kathy's age = K
Cat's age = C
.
K = C +6
.
Next year...
K+1
C+1
K+1 = 2(C+1)
.
Substitute K =C+6
.
C+6 +1 = 2(C+1)
C+7 = 2C +2
-C = -5
C = 5
.
The cat is 5 years old.
so
Kathy is 11 years old (5+6=11).
.
Next year, Kathy will be 12 and the cat will be 6, so Kathy will be twice the cat's age.
.
Done.